OpenStudy (anonymous):

A father wishes to provide P40,000 for his son on his son’s 21st birthday. How much should he deposit every 6 months in a savings account which pays 3% compounded semi-annually, if the first deposit was made when the son was 3 ½ years old?

OpenStudy (anonymous):

@mathstudent55

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@goformit100

OpenStudy (mathstudent55):

The interest rate is 3% per year compounded 2 times per year. The time is 21 - 3.5 years = 17.5 years The future amount is P40,000 We need the formula for compound interest, and then to solve it for present value.

OpenStudy (mathstudent55):

\(\large F = P \left(1 + \dfrac{r}{n}\right)^{nt} \) where F = future value P = present value r = annual interest rate written as a decimal n = number of compounding periods per year t = number of years

OpenStudy (mathstudent55):

Plug in all your numbers in the formula, and solve for P, the present value.

OpenStudy (anonymous):

but its a deffred annuity @mathstudent55

OpenStudy (mathstudent55):

Sorry. You are correct. That is the formula if you made a single deposit now (present value), and you want a future value later. For an annuity, there is a different formula. \(F = A\times \dfrac{n(1 + \frac{r}{n})^{nt} - n }{r } \) where r = decimal interest rate n = number of compoundings per year t = number of years A = periodic payment F = future value

OpenStudy (anonymous):

i got the answer thanks guys :)